OpenStudy (anonymous):

Find the fourth-degree polynomial function with zeros 2, -2, 2i, and -2i. Write the function in factored form.

OpenStudy (faiqraees):

x= root (x-root) is a factor of equation factor*factor*factor*factor = fourth degree polynomial

OpenStudy (anonymous):

I don't understand. @FaiqRaees

OpenStudy (danjs):

recall when you solved something like x^2 + 3x + 2 = 0 (x + 1)(x + 2) = 0 you factored like that, and then set each quantity =0 to solve for x

OpenStudy (danjs):

do that backwards here, they gave you the x values for this thing, make quantities that will equal zero with those x values... like (x - 2i) +2i will make that zero

OpenStudy (anonymous):

So, it will be something like: x(x+2)(x-2)(x+2i)(x-2i)? sorry, i'm new at this and really bad at it

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

yes good, but the x in front says that x=0 would be a root,

OpenStudy (anonymous):

Alright. So no x in front.

OpenStudy (danjs):

(x+2)(x-2)(x+2i)(x-2i) = 0 that is good enough for me, they prolly want it in ax^4 + bx^3 + cx^2 + dx + e = 0 so you would have to multiply it all out

OpenStudy (danjs):

(x^2 - 4) ( x^2 - 4i^2) = 0 (x^2 - 4) ( x^2 - 4*(-1)) = 0 (x^2 - 4) ( x^2 + 4) = 0

OpenStudy (anonymous):

It was right! Wow. That was a lot easier then I thought.

OpenStudy (anonymous):

Thank you so much!